OpenStudy (anonymous):

Show that 2^n > n^2 for all integers n > 4. Basis: 2^(5) > 5^2 which is true I.hypothesis: 2^k > (k)^2 I. step: 2^(k+1) > (k+1)^2 (2^k)*2 > (k+1)(k+1) (2^k)*2 > k^2+2k+1 I'm stuck here. I can't find a way to reduce this, either by relating it to the IH or find a "middle inequality" to prove the induction. Any tips to help me further the question?

OpenStudy (usukidoll):

we need the k step and the k+1 step... so I usually expand the right as a hint for me to figure out what I need to do with the left

OpenStudy (anonymous):

which I did haha. Im stuck at it, thats the thing. If 2^k * 2 > k^2 +2k + 1. There's no way for me to show that 2> 2k + 1 because that would be false in itself. However, when grouped with the 2^k and k^2, it seems that its just unknown

OpenStudy (usukidoll):

wait a sec.. this could be a different version of the induction... like we do the same find k and k+1 but I think you have to split the 2^(k+1) = 2^k 2^1 = 2(2^k) and then subsitute what 2^k is or what k is crap I need to do this again .... maybe it's in my stackexchange account

OpenStudy (anonymous):

i would post on stackexchange but people there are... wingspans a lot of the time

OpenStudy (usukidoll):

oh yeah...

OpenStudy (usukidoll):

I remember having to substitute for that 2^k

OpenStudy (anonymous):

I can't figure it out still, I got to separating the expression but after that it gets rocky

OpenStudy (usukidoll):

because that left side of yours is wrong. 2^{k+1} should be 2^k (2)

OpenStudy (usukidoll):

nughhh even I'm stuck...

OpenStudy (usukidoll):

I did this problem before... wait let me check my book

OpenStudy (usukidoll):

wow there's an example sheet that I did a year ago.. I can scan for you if you want

OpenStudy (anonymous):

that would be awesome if you dont mind haha

OpenStudy (anonymous):

thatd be even too nice o.o

OpenStudy (usukidoll):

OH! I GET IT NOW! 2k^2 = k^2 + k^2 and then you use pascal's triangle

OpenStudy (usukidoll):

k I'll scan it

OpenStudy (usukidoll):

first k^2 is left alone and then you subsitute the second k^2 with 2k and then add 1

OpenStudy (anonymous):

wtf is pascale triangle. and wth LOL Im lost you lost me 100% there

OpenStudy (usukidoll):

2^k = k^2 you use subsitutuion then 22^k = 2k^2 = k^2 +k^2 skipped a step

OpenStudy (usukidoll):

2^k = k^2 from that k hypothesis

OpenStudy (usukidoll):

so you had 2(2^k) using the k hypothesis you subsitute 2^k with k^2 leaving 2(k^2) = 2k^2 which is actually 2k^2 = k^2 +k^2 and then you use pascal's triangle on the second k^2 and add 1 so that's k^2+2k+1 which I got left = right

OpenStudy (anonymous):

Ive never learned pascale triangle hahaha, ill check it out. But how did you go from 2(2^k) --> 2(k^2) if 2^k > k^2 at the start ?

OpenStudy (usukidoll):

what I did a year ago is similar to yours only bigger numbers

OpenStudy (usukidoll):

o-o 2^k > k^2 was your k hypothesis right?

OpenStudy (usukidoll):

you subsitutue 2^k with k^2 so you have (2^1)(k^2) = 2k^2 and then you split that into k^2+k^2

OpenStudy (usukidoll):

if you don't use substitution... you will be stuck and never get it done.. even I did it on my paper check the first problem with the 4

OpenStudy (usukidoll):

do you see 4^k > k^4 ? I substituted it

OpenStudy (anonymous):

nono , I'm not denying what you're saying haha. but if your k hypothesis is 2^k > k^2. How can 2^(K+1) = 2(k^2) if 2^k > k^2

OpenStudy (usukidoll):

x.x I used to get stuck on that... my adviser told me to subsitute this is a loop question isn't it... first of all the left side of your k+1 is wrong

OpenStudy (usukidoll):

2^(k+1) = 2^k(2^1) = 2^k(2)

OpenStudy (usukidoll):

sorry I froze.

OpenStudy (anonymous):

how is my k+1 wrong O.o

OpenStudy (usukidoll):

from the k hypothesis you have 2^k > k^2 so you need to exchange 2^k with k^2 ... the way you split it up is wrong

OpenStudy (usukidoll):

it's similar to exponentials... e^(x+1) is the same as e^x(e^1)

OpenStudy (usukidoll):

so 2^(k+1) is really 2^k(2^1) = 2^k(2) and I did that too on my paper

OpenStudy (anonymous):

thats what I said O.o 2^(k+1) is the same as (2^k)(2) haha

OpenStudy (usukidoll):

not in the original post when you asked the question

OpenStudy (anonymous):

whoops tryping error then! :) but you then said 3k * 3 = 3k^3 right?

OpenStudy (anonymous):

in your paper that is

OpenStudy (usukidoll):

O_____________O! I'm focusing on the first question

OpenStudy (anonymous):

WHOOPS wrong one lol sorry

OpenStudy (usukidoll):

when i splitted it up again I exchanged 4^k with k^4 using the k hypothesis leave you with 4k^4 that is actually = k^4+k^4+k^4+k^4 then you add 1 at the end and use pascal's triangle to figure out the 2-4th k^4s except yours should be easy

OpenStudy (anonymous):

same idea though you said 4^k * (4) is equivalent to 4k^4

OpenStudy (anonymous):

I mean you exchanged an exponent with a base, to me thats defying the law of math o.o

OpenStudy (usukidoll):

tell that to my adviser. he taught me how to do it

OpenStudy (anonymous):

LOL im sorry its just weird to me, i didnt know you could do it. Im not trying to say anything bad, im far from having a phd in math or anything. It just seems... odd that 4^k * (4) = 4k^4

OpenStudy (usukidoll):

UGH! YOu just exchange 4^k with k^4 using the k hypothesis!!!

OpenStudy (anonymous):

Its okay I just verified it using numbers, its true, but its just weird lol cause we have an inequality. if 4^k = k^4 in k I would do it but we have greater then, hence why I felt weird about doing it. thats all haha